# A Chinese Prompt Attack Dataset for LLMs with Evil Content

### v1

@@ Conclusion
In this paper, we introduce CPAD, which is a Chinese prompt attack dataset containing 10050 samples. We conclude three dimensions, content, template and goal from the perspective of attackers, which is different from previous studies. We utilize ChatGPT to extend manually-written seed samples, and only keep the successful prompts against three popular Chinese LLMs, where the evaluation prompts are constructed given the attacking goals and contents. We conduct analysis on responses from another four LLMs, including ChatGPT. The evaluation shows that CPAD has an attack success rate of around 70% against the LLMs. We also fine-tune Baichuan-13B-Chat using parts of CPAD, which improves the safety significantly.
Our analysis reveals the weakness of LLMs including ChatGPT, and indicates that there is still significant room for improvement in terms of safety. CPAD may contribute to further prompt attack studies.
Appendix A Fine-Tuning Implementation Details
We adopt Baichuan-13B-Chat as the LLM, which is the aligned version in the Baichuan-13B series of models. The checkpoint is downloaded from Huggingface. We adopt LoRA for parameter-efficient fine-tuning.

### v2

@@ Conclusion
In this paper, we introduce a pipeline to construct high-quality prompt attack samples, along with a Chinese prompt attack dataset called CPAD containing 10050 samples. There are three key dimensions, content, template and goal from the perspective of attackers, which is different from previous studies. We utilize GPT-3.5 to extend manually-written seed samples, and only keep the successful prompts against three popular Chinese LLMs, where the evaluation prompts are constructed given the attacking goals and contents. We conduct analysis on responses from another four LLMs, including GPT-3.5. The evaluation shows that CPAD has an successful-attacking rate of around 70% against the LLMs. We also fine-tune Baichuan-13B-Chat using parts of CPAD, which improves the safety significantly.
Our analysis reveals the weakness of LLMs including GPT-3.5, and indicates that there is still significant room for improvement in terms of safety. CPAD may contribute to further prompt attack studies.